# Continuous-Time State Estimation Methods in Robotics: A Survey

### v1

@@ IV-C5 Other Approaches
Le Gentil et al. [210] propose a novel method for event camera motion compensation using GPs. By treating camera motion as motion in the image frame over short time periods, they model each DoFs as independent zero-mean GPs with Gaussian kernels. The inducing values (i.e., set of observations) of these are then the hyperparameters of another GP representing a continuous occupancy field of events in the image plane. By optimizing these through log-marginal likelihood maximization, this occupancy field, and hence a continuous distance field as its negative logarithm, can be inferred and used for homography registration and pattern tracking. Le Gentil and Vidal Calleja [211] use a GP pose representation in the extrinsic calibration of an IMU with pose measurement reference frame (e.g., motion capture or robotic arm frames). Kapushev et al. [199] propose a SLAM algorithm evaluated in 2D, which models the pose as a GP with a Random Fourier Features (RFF) approximation of the Gaussian kernel and uses either a discretized motion model or smoothing splines for the prior mean function.
@@ V-A1 Knot Selection
The optimal selection of knots remains one of the most important unresolved problems of the spline method. The simplest and most common approach is to use uniform splines, tuning the frequency according to the application. While uniform splines benefit from a fixed blending matrix that needs to be computed only once, they implicitly assume that excitation of the state evolution (e.g., robot motion) remains constant. Hence, uniform knots can lead to underfitting or overfitting of the trajectory in settings with varying dynamics [31]. Furthermore, a control point is only well-constrained if there are sufficient measurements on the associated segments, meaning that optimization can be poorly posed if the knot frequency is too high relative to the measurement frequency.
@@
Several alternative schemes have been offered. Dubé et al. [101] propose several distributions for the knots within the sliding window of the optimization. These include an exponential distribution, a frequency analysis distribution based on average correction signal power, and a uniform distribution, potentially in combination. Hug and Chli [31] discuss optimization of the knots themselves and derive Jacobians for them. However, this idea has not yet been pursued experimentally. Regularizing motion terms [20, 45, 22] may avoid the issue of under-constrained control points; however, it may add assumptions conflicting with the implicit spline constraints.
@@ V-A2 Certifiability & Recovery Guarantees
Important open research questions arise because the optimization methods commonly used for estimation may converge to bad local minima or fail to converge altogether. Indeed, depending on the measurement modalities used, the residuals of the factor graph (i.e., optimization cost terms) may be non-convex. Particularly, if states are confined to manifold spaces, this may lead to non-convex constraints and exacerbate the problem. Multiple recent works [219, 220, 221] show that non-convexity may give rise to bad local minima in which local solvers such as GN can get stuck. Estimating splines may also suffer from this, although the inherent regularizing properties could help escape local minima.
@@
V-A3 Incremental Optimization
Sliding-window optimization has been used to facilitate real-time operation [99, 100, 101, 104, 54, 165, 55, 168, 170, 171, 174, 175]. Other incremental schemes, such as those based on the Bayes tree [218], have not been explored. Various graphical models have been presented for splines [164, 165, 55], however only very recent works [168, 170, 171, 175] have made serious efforts to represent spline estimation with factor graphs and use these to design marginalization strategies. Leveraging these to incorporate incremental optimizers like iSAM2 [218] could be explored in future work.
V-A4 Initialization
Initialization of control points is rarely discussed in the literature. However, appropriate initialization certainly affects convergence time and optimality. Simple strategies for initializing new control points include using the value of the previous control point, or IMU dead reckoning. However, the state estimate may be far from the control point, especially for highly dynamic motion and high-order splines, and no inverse interpolation function exists to compute control points from the state. One strategy adopted by Huang et al. [163] was to use a spline approximation algorithm from Piegl and Tiller [222] for initialization. Li et al. [177] devise a sensor-specific coarse-to-fine procedure to initialize control points, biases, and extrinsic calibration parameters. Kang and Park [129] offer initialization equations for different parameterizations of orientation splines based on boundary conditions. Yet, initialization remains an open problem.
V-A5 Covariance Interpolation
It is often desirable to infer the uncertainty of the state at the interpolation time from the uncertainty of the control points. In fact, the covariance can be interpolated for vector-valued states, with some limited empirical results demonstrated [20, Figure 6]. Bibby and Reid [131] provide a derivation for their non-uniform cubic splines, which they applied to dynamic SLAM in 2D, using the Jacobian of the spline interpolation function. State covariance is also estimated by the probabilistic filter proposed by Li [172]. However, no works have provided a covariance interpolation formulation for Lie group splines.
@@ V-B2 Certifiability & Recovery Guarantees
In contrast with the spline method, GP motion priors can be effortlessly integrated with this paradigm, as shown recently for the examples of CT range-only localization [224] and pose-graph optimization [215]. Indeed, to incorporate a GP motion prior, the state is augmented with the necessary number of temporal derivatives, after which the motion prior regularization always takes the form of additional inter-state quadratic cost terms, which can be readily included in the QCQP formulation. However, augmenting the state and regularizing the problem in this way may affect the tightness of the SDP relaxations of the QCQP, which, in turn, can break certifiability. For pose-graph optimization in particular, a significant number of so-called redundant constraints are required for tightness at reasonable noise levels when using the WNOA prior [215, Equation (57)], which breaks typical constraint qualifications and significantly increases the computational complexity of the used SDP solvers. Measures to increase and (a-priori) understand tightness and to speed up SDP solvers are crucial for advancing the practicality of these methods.
@@
V-B3 Complex Process Models
Tang et al. [62] show that selecting a process model prior that does not represent state evolution leads to biased estimation. Thus, it is clear that a prior that most appropriately reflects the state evolution should be chosen. The TGP method inherently allows for complex process dynamics and control inputs to regularize the trajectory and guide interpolation. However, thus far, only relatively simple models, such as the WNOA or WNOJ priors, have been examined. While they can offer computational savings and require tuning only a small number of hyperparameters, they may not accurately reflect the process dynamics of complex robotics systems or even simple processes from other fields. Questions about how beneficial and computationally tractable more complex priors would be and how feasible these would be for real-time applications, remain open for this method.
Some recent works have used learning in the modeling of system dynamics for fast online use [225, 226], and it would be reasonable to consider this idea in the computation or augmentation of the prior. A principled method for tuning hyperparameters (e.g., ) is the maximization of the marginal log-likelihood using a training data set [35] (over hand-tuning [62, 58, 216] or fitting to data [59]), such as with the Singer prior [198]. While this typically requires ground truth data, the EM parameter learning framework extending ESGVI [196] proposed by Wong et al. [197] demonstrates that hyperparameters can be learned from the original noisy measurement data. The Singer prior [198] is an example of an alternative approach to process modelling, whereby an SDE is derived from factorization of the spectral density of a GP prior, such as a kernel from the Matérn family [227, 228]. It remains to be seen what methods could be effectively employed to learn reasonable GP priors for complex processes.
@@ VI-A Temporal Splines
Moreover, splines whose basis functions form a partition of unity (i.e., they are nonnegative, and their sum is one) exhibit the convex hull property, which guarantees that each spline segment is wholly contained within the convex hull of its control points. This is leveraged by many trajectory planning works [236, 237, 238, 239, 240, 241, 242], since it substantially simplifies obstacle avoidance constraints [243]. B-splines are the typical choice of splines for trajectory planning, as they can guarantee smoothness and high continuity by construction without imposing explicit constraints. However, the convex hull of the B-Spline control points is typically much larger than the segment itself. To reduce conservativeness, some works [244] use B-spline control points as decision variables in trajectory planning. However, the Bézier control points of each interval of this B-spline are used in the obstacle avoidance constraints. The MINVO control points [245, 246] allow for even tighter enclosures and have been leveraged for obstacle avoidance [246].
@@ VI-B Temporal Gaussian Processes (TGPs)
GPs have been deployed for a large number of other tasks in robotics. Zhang and Scaramuzza [264] demonstrates the capacity for temporal GPs with Gaussian kernels to represent ground truth trajectories, proposing generalized absolute and relative errors for principled trajectory evaluation. GPs with Gaussian kernels have also been used in Bayesian filters, such as for replacing or enhancing the parametric prediction and observation models [265, 44, 266]. The use of GPs as surrogate observation models has been explored for signal-strength (e.g., WiFi, magnetic flux) state estimation applications [267, 268, 269, 270], and as dynamics prediction models within the control community [271, 272, 273, 274, 275]. In informative path planning applications, GPs have been used for representing scalar environmental fields, like temperature, salinity, methane concentration, magnetic field intensity, and terrain height [276, 277, 278, 279, 280, 281]. GPs have also seen extensive use in mapping, including occupancy maps [282, 283, 284, 285, 286, 287, 288, 289, 290, 291, 292, 293], implicit surfaces [294, 295, 296, 297, 298, 299, 300, 301, 302, 303, 304, 305], their combination [306], gradient maps [307, 308, 309, 310] and distance fields [302, 305, 311, 210, 312, 313]. They have also been used as a representation for local perception and exploration uncertainty [314]. However, these works did not use LTV-SDE priors to obtain exact sparsity in GP regression.777Although, some works [306] attain sparsity in the covariance matrix by using a piecewise kernel that is zero beyond a certain distance [315].
Mukadam et al. [316] were the first to apply exactly sparse TGPs to the problem of motion planning with the GPMP algorithm, in particular for a 7-DoF robotic arm, addressing the large number of states needed by discrete-time planners. This was extended to factor graph optimization, in GPMP2 [317] (later used by Maric et al. [191]), GPMP-GRAPH [318], and iGPMP2 [317, 58], before unification with state estimation in STEAP [188, 189], and learning from demonstration in CLAMP [190]. CLAMP is unique from other works in that the transition matrix and bias are learned for particular tasks from trajectory demonstrations with linear ridge regression. GPMP2 was also extended to dGPMP2 [319], which through a self-supervised end-to-end training framework, allowed the learning of factor covariance parameters, such as , an idea which could be applied to state estimation in future work. Inspired by these works, Cheng et al. [320] adopts an exactly sparse GP for path planning in an autonomous driving context, parameterizing their WNOJ prior by arc length instead of time. Lilge et al. [321] similarly parameterizes by arc length to model the shape of a continuum robot using a white-noise-on-strain-rate (analogous to WNOA) LTV-SDE prior derived from the Cosserat rod model, with state consisting of pose and generalized strain.
@@ VII Conclusion
| / | / | Dynamic Object Poses/Velocities |

### v2

@@ IV-C5 Other Approaches
Shen et al. [46] adapt the TGP method to an EKF for the first time to achieve real-time dropout-resilient multi-LiDAR odometry. They create point-to-voxel residuals with a degeneracy-aware point sampling strategy and an incremental, probabilistic voxel-based map. Le Gentil et al. [221] propose a novel method for event camera motion compensation using GPs. By treating camera motion as motion in the image frame over short time periods, they model each DoF as independent zero-mean GPs with Gaussian kernels. The inducing values (i.e., set of observations) of these are then the hyperparameters of another GP representing a continuous occupancy field of events in the image plane. By optimizing these through log-marginal likelihood maximization, this occupancy field, and hence a continuous distance field as its negative logarithm, can be inferred and used for homography registration and pattern tracking. Le Gentil and Vidal Calleja [222] use a GP pose representation in the extrinsic calibration of an IMU with a pose measurement reference frame (e.g., motion capture or robotic arm frames). Kapushev et al. [209] propose a SLAM algorithm evaluated in 2D which models the pose as a GP, with a Random Fourier Features (RFF) approximation of the Gaussian kernel and a discretized motion model or smoothing splines for the prior mean.
@@ V-A1 Knot Selection
The optimal selection of knots remains an important unresolved problems of the spline method. The simplest and most common approach is to use uniform splines, tuning the frequency according to the application. While uniform splines benefit from a fixed blending matrix that needs to be computed only once, they implicitly assume that excitation of the state evolution (e.g., robot motion) remains constant. Hence, uniform knots can lead to underfitting or overfitting of the trajectory in settings with varying dynamics [31]. Furthermore, a control point is only well-constrained if there are sufficient measurements on the associated segments, meaning that optimization can be poorly posed if the knot frequency is too high relative to the measurement frequency.
@@
Several other schemes have been offered. Dubé et al. [104] propose several distributions for the knots within the sliding window of the optimization. These include an exponential distribution, a frequency analysis distribution based on average correction signal power, and a uniform distribution, potentially in combination. Hug and Chli [31] discuss optimization of the knots themselves and derive Jacobians for them. However, this idea has not yet been pursued experimentally. Regularizing motion terms [20, 47, 22] may avoid the issue of under-constrained control points; however, it may add assumptions conflicting with the implicit spline constraints.
@@ V-A2 Certifiability & Recovery Guarantees
Important open research questions arise because the optimization methods commonly used for estimation may converge to bad local minima or fail to converge altogether. Indeed, depending on the measurement modalities used, the residuals of the factor graph (i.e., optimization cost terms) may be non-convex. Particularly, if states are confined to manifold spaces, this may lead to non-convex constraints and exacerbate the problem. Multiple recent works [231, 232, 233, 234] show that non-convexity may give rise to bad local minima in which local solvers such as GN can get stuck. In the face of these convergence issues, the global minimum can still be recovered from a tight but more costly convex relaxation (e.g., for initialization [232]).
@@
V-A3 Complex Processes
The piecewise-polynomial interpolation of splines may be problematic in estimating high-frequency or hybrid dynamic systems, such as those that may occur in the presence of vibration or impacts. Rather than increase knot frequency or spline order, which raises the computational load, Li et al. [187] propose using the RMSE of the accelerometer measurements to adaptively select between using IMU measurements directly or preintegrating them as a form of low-pass filter. Additionally, they apply constant-velocity factors to the roll and pitch when the acceleration is sufficiently low. Resilience to non-polynomial and hybrid dynamics remains a challenging problem for the spline method.
V-A4 Incremental Optimization
Sliding-window optimization can facilitate real-time operation [102, 103, 104, 107, 56, 172, 57, 175, 177, 178, 182, 183]. Other incremental schemes, such as those based on the Bayes tree [230], have not been explored. Various graphical models have been presented for splines [171, 172, 57], however only very recent works [175, 177, 178, 183] have made serious efforts to represent spline estimation with factor graphs and use these to design marginalization strategies. Leveraging these to incorporate incremental optimizers like iSAM2 [230] could be explored in future work.
V-A5 Initialization
Initialization of control points is rarely discussed in the literature. However, appropriate initialization certainly affects convergence time and optimality. Simple strategies for initializing new control points include using the value of the previous control point, or IMU dead reckoning. However, the state estimate may be far from the control point, especially for highly dynamic motion and high-order splines, and no inverse interpolation function exists to compute control points from the state. One strategy adopted by Huang et al. [170] was to use a spline approximation algorithm from Piegl and Tiller [235] for initialization. Li et al. [185] devise a sensor-specific coarse-to-fine procedure to initialize control points, biases, and extrinsic calibration parameters. Kang and Park [136] offer initialization equations for different parameterizations of orientation splines based on boundary conditions.
V-A6 Covariance Interpolation
It is often desirable to infer the uncertainty of the state at the interpolation time from the uncertainty of the control points. To a limited extent, covariance interpolation for vector-valued splines has been demonstrated [20, Figure 6][138]. Bibby and Reid [138] provide a derivation for their non-uniform cubic splines, applied to dynamic SLAM in 2D, using the Jacobian of the spline interpolation function. State covariance is also estimated by the probabilistic filter proposed by Li [45]. However, no works provide a covariance interpolation formulation for Lie group splines.
@@ V-B2 Certifiability & Recovery Guarantees
In contrast with the spline method, GP motion priors can be effortlessly integrated with this paradigm, as shown recently for the examples of CT range-only localization [237] and pose-graph optimization [226]. Indeed, to incorporate a GP motion prior, the state is augmented with the necessary number of temporal derivatives, after which the motion prior regularization always takes the form of additional inter-state quadratic cost terms, which can be readily included in the QCQP formulation. However, augmenting the state and regularizing the problem in this way may affect the tightness of the SDP relaxations of the QCQP, which, in turn, can break certifiability. For pose-graph optimization in particular, a significant number of so-called redundant constraints are required for tightness at reasonable noise levels when using the WNOA prior [226, Equation (57)], which breaks typical constraint qualifications and significantly increases the computational complexity of the used SDP solvers. Measures to increase and (a-priori) understand tightness and to speed up SDP solvers are crucial for advancing the practicality of these methods.
@@
V-B3 Complex Processes
Tang et al. [64] show that selecting a process model prior that does not represent state evolution leads to biased estimation. Thus, it is clear that a prior that most appropriately reflects the state evolution should be chosen. TGPs inherently allows for complex process dynamics and control inputs to regularize the trajectory and guide interpolation. However, only relatively simple models, such as the WNOA or WNOJ priors, have been examined thus far. While efficient to compute and require tuning only a small number of hyperparameters, they do not accurately approximate the dynamics of all systems. Systems with hybrid dynamics or impacts (e.g., legged robotics) may be particularly problematic since impulse-like accelerations induce perceived velocity discontinuities, violating the assumptions of existing priors. Further research is needed into the feasibility and tractability of system-specific priors and control input utilization.
Some recent works have used learning in the modeling of system dynamics for fast online use [238, 239], and it would be reasonable to consider this idea in the computation or augmentation of the prior. A principled method for tuning hyperparameters (e.g., ) is the maximization of the marginal log-likelihood using a training data set [35] (over hand-tuning [64, 60, 227] or fitting to data [61]), such as with the Singer prior [207]. While this typically requires ground truth data, the EM parameter learning framework extending ESGVI [205] proposed by Wong et al. [206] demonstrates that hyperparameters can be learned from the original noisy measurement data. The Singer prior [207] is an example of an alternative approach to process modeling, whereby an SDE is derived from factorization of the spectral density of a GP prior, such as a kernel from the Matérn family [240, 241]. It remains to be seen what methods could be effectively employed to learn reasonable GP priors for complex processes.
V-C Evaluation and Benchmarking
Public benchmarks and appropriate metrics must be established with which CT algorithms can be compared and evaluated. While many works provide comparisons on public benchmarks to other hand-picked methods [177, 182, 30, 223, 178, 76, 180], they are frequently inconsistent, adding ambiguity. Since CT methods can fuse high-rate asynchronous measurements, benchmarks must consider how to provide sensor data and ground truth references. For example, MCD [186] recently leveraged survey-grade maps to optimize a “ground truth” B-spline from LiDAR-inertial data. Existing algorithms are usually compared with aggregated absolute or relative trajectory errors, which are insufficient metrics for understanding their complete behavior. Zhang and Scaramuzza [242] demonstrate the capacity for temporal GPs with Gaussian kernels to represent ground truth trajectories and propose generalized absolute and relative errors for principled trajectory evaluation.
@@ VI-A Temporal Splines
Moreover, splines whose basis functions form a partition of unity (i.e., nonnegative, summing to one) exhibit the convex hull property, which guarantees that each spline segment is wholly contained within the convex hull of its control points. This is leveraged by many trajectory planning works [250, 251, 252, 253, 254, 255, 256], since it substantially simplifies obstacle avoidance constraints [257]. B-splines are the typical choice of splines for trajectory planning, as they can guarantee smoothness and high continuity by construction without imposing explicit constraints. However, the convex hull of the B-Spline control points is typically much larger than the segment itself. To reduce conservativeness, some works [258] use B-spline control points as decision variables in trajectory planning, but then, the Bézier control points of each interval of this B-spline are used in the obstacle avoidance constraints. The MINVO control points [259, 260] allow for even tighter enclosures and have been leveraged for obstacle avoidance [260].
@@ VI-B Temporal Gaussian Processes (TGPs)
GPs have been deployed for a large number of other tasks in robotics. GPs with Gaussian kernels have often been used in Bayesian filters, such as for replacing or enhancing the parametric prediction and observation models [278, 44, 279]. The use of GPs as surrogate observation models has been explored for signal-strength (e.g., WiFi, magnetic flux) state estimation applications [280, 281, 282, 283], and as dynamics prediction models within the control community [284, 285, 286, 287, 288]. In informative path planning applications, GPs have been used for representing scalar environmental fields, like temperature, salinity, methane concentration, magnetic field intensity, and terrain height [289, 290, 291, 292, 293, 294]. GPs have also seen extensive use in mapping, including occupancy maps [295, 296, 297, 298, 299, 300, 301, 302, 303, 304, 305, 306], implicit surfaces [307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 318], their combination [319], gradient maps [320, 321, 322, 323] and distance fields [315, 318, 324, 221, 325, 326]. They have also been used as a representation for local perception and exploration uncertainty [327]. However, these works did not use LTV-SDE priors to obtain exact sparsity in GP regression.777Although, some works [319] attain sparsity in the covariance matrix by using a piecewise kernel that is zero beyond a certain distance [328].
Mukadam et al. [329] were the first to apply exactly sparse TGPs to the problem of motion planning with the GPMP algorithm, in particular for a 7-DoF robotic arm, addressing the large number of states needed by discrete-time planners. This was extended to factor graph optimization, in GPMP2 [330] (later used by Maric et al. [200]), GPMP-GRAPH [331], and iGPMP2 [330, 60], before unification with state estimation in STEAP [197, 198], and learning from demonstration in CLAMP [199]. CLAMP is unique from other works in that the transition matrix and bias are learned for particular tasks from trajectory demonstrations with linear ridge regression. GPMP2 was also extended to dGPMP2 [332], which through a self-supervised end-to-end training framework, allowed the learning of factor covariance parameters, such as , an idea which could be applied to state estimation in future work. Inspired by these works, Cheng et al. [333] adopt an exactly sparse GP for path planning in an autonomous driving context, parameterizing their WNOJ prior by arc length instead of time. Lilge et al. [334, 335] similarly parameterize by arc length to model the shape of a continuum robot using a white-noise-on-strain-rate (analogous to WNOA) LTV-SDE prior derived from the Cosserat rod model, with state consisting of pose and generalized strain. They recently explore including both temporal and spatial domains jointly [336].
@@ VII Conclusion
| // | // | Dynamic Object Pose/Velocity/Acceleration |
| // | // | Dynamic Object Position/Velocity/Acceleration |